# Safe and Generalized end-to-end Autonomous Driving System with Reinforcement Learning and Demonstrations

### 23 Jan 2024 v2

@@ 1 Introduction
In conclusion, the existing IL and RL methods can not well solve the safety, sampling efficiency and generalization of autonomous vehicles. To solve above problems, in this paper, we introduce a Safe and Generalized end-to-end Autonomous Driving System (SGADS) for complex and various scenarios. Our SGADS incorporates variational inference with normalizing flows, enabling the intelligent vehicle to accurately predict future driving trajectories. Moreover, we design safety constraint for intelligent vehicle. We further propose a behavior cloning as to augment search process of agent. We also contribute a demonstration dataset by having human experts drive using the G29 steering wheel. The experimental result shows our SGADS greatly augments safety performance, exhibits strong generalization in multiple testing maps, and enhances the training efficiency of intelligent vehicles in intricate urban scenarios compared to the existing methods. Moreover, we designed and explored up to 12 different data input types to evalutate their impact on the safety performance of intelligent vehicles. To sum up, our contributions are:
@@ 4.1.2 Measure Safety
Safety is defined as the conditions in which an autonomous vehicle in its environment can follow a predetermined route without colliding with surrounding vehicles, avoid major deviations from a predetermined trajectory, avoid stopping on the road surface, avoid contact with a roadside wall or guardrail, and successfully navigate a complex intersection or roundabout without a collision. In this experimental environment setting, if the intelligent vehicle fails to comply with the above regulatory constraints, it will be penalized, resulting in the termination of the current episode. In addition, when an episode ends, all scenes and vehicles are forcibly reset, and the reset of the intelligent car is also considered as the vehicle experienced a collision, resulting in the end of the episode. Therefore, we measure the safety using the Average Driving Distance(Avg Dis), which you can think of as the average distance a self-driving car can safely travel in multiple episodes. When intelligent vehicle travels a longer driving distance within the episode, it indicates that its safety performance is higher. Simultaneously, as the episode length increases and our testing extends, the vehicle is more likely to have a longer safe driving distance. Additionally, we evaluate the ability to safely travel the maximum distance over multiple episodes using Max Driving Distance(Max Dis), calculated as the maximum value of the driving distance across multiple episodes.

### 24 Jan 2024 v3

@@ 1 Introduction
In conclusion, the existing IL and RL methods can not well solve the safety, sampling efficiency, generalization and prediction for future driving trajectories in autonomous driving. To solve above problems, in this paper, we introduce a Safe and Generalized end-to-end Autonomous Driving System (SGADS) for complex and various scenarios. Our SGADS incorporates variational inference with normalizing flows, enabling the intelligent vehicle to accurately predict driving trajectories. Moreover, we design safety constraint for intelligent vehicle. We further propose a behavior cloning as to augment search process of agent. We also contribute a demonstration dataset by having human experts drive using the G29 steering wheel. The experimental result shows our SGADS greatly augments safety performance, exhibits strong generalization in multiple testing maps, and enhances the training efficiency of intelligent vehicles in intricate urban scenarios compared to the existing methods. Moreover, we designed and explored up to 12 different data input types to evalutate their impact on the safety performance of intelligent vehicles. To sum up, our contributions are:
@@ 4.1.2 Measure Safety
In this experimental environment setting, high safety is defined as the conditions in which an intelligent vehicle in its environment can follow a randomly predetermined route without colliding with surrounding vehicles, avoid major deviations from a randomly predetermined trajectory, avoid stopping on the road, avoid colliding with a roadside wall or guardrail, and successfully navigate a complex intersection or roundabout without any collision. If the intelligent vehicle fails to comply with any of the above regulatory constraints, it will be penalized, resulting in the termination of the current episode in our setting. In addition, when an episode ends, all scenes and vehicles are forcibly and randomly reset, resulting in the termination of the episode. Therefore, we use the Average Driving Distance (Avg Dis) to measure safety, indicating the average distance that an intelligent vehicle can safely drive in multiple episodes. When the intelligent vehicle drives a longer driving distance within the episode, it indicates that its safety performance is higher. Simultaneously, as the episode length increases and our testing extends, the intelligent vehicle is more likely to have a longer safe driving distance. Additionally, we evaluate the ability to safely travel the maximum distance over multiple episodes using Max Driving Distance(Max Dis), calculated as the maximum value of the driving distance across multiple episodes.